# From Cash to Cashless: UPI’s Impact on Spending Behavior among Indian Users

### v1

@@ Abstract.
The emergence of digital payment systems has transformed how individuals conduct financial transactions, offering convenience, security, and efficiency. One groundbreaking innovation making waves in the Indian financial landscape is the Unified Payments Interface (UPI), developed by the National Payments Corporation of India (NPCI). Existing work has explored how digital payments benefit a country’s economy and GDP. However, our study explores how the introduction of UPI has influenced spending behavior among Indian users on an ”individual” level. We gathered 235 valid survey responses encompassing diverse demographics and conducted semi-structured interviews with 20 survey respondents. Approximately 75% of the survey respondents reported increased spending due to UPI, with only 7% indicating reduced spending. Significantly, 91.5% of the respondents reported satisfaction with their UPI usage. Also 95.2% of the survey respondents found making payments via UPI convenient. Our research also provides suggestions for UPI applications and various stakeholders to enhance digital payment systems, enabling users to make informed decisions and fostering responsible financial management.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM Conference on Computer-Supported Cooperative Work and Social Computing; November 09–13, 2024; San Jose, Costa Ricabooktitle: CSCW’24: ACM Conference on Computer-Supported Cooperative Work and Social Computing, November 09–13, 2024, San Jose, Costa Ricaprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
The term ”E-payment” is comprehensive, illustrating diverse proportions of E-payments through different channels. Despite attempts to define electronic payment, a clear, universally accepted definition is still lacking. An E-payment is generally understood as an economic situation where the buying and selling of goods are handled without physical cash, often facilitated through E-payment or checks (Pau, 2012). Digital payments have become integral to modern financial landscapes, offering swift, secure, and convenient electronic transactions that transcend traditional methods. This global shift towards digital transactions has been accelerated by the widespread adoption of internet and mobile technologies, reshaping financial practices across the world (Alkhowaiter, 2020). In 2022 alone, the volume of real-time digital transactions reached an unprecedented 195 million (89., 2023). Notably, 64 % of adults globally engaged in at least one digital payment transaction, with high-income economies leading at 95% (glo, 2021).
@@
India, in particular, has emerged as a global leader in digital transactions, accounting for 46 % of the total global digital transactions in 2022 (89., 2023). At the forefront of this digital revolution is the Unified Payments Interface (UPI), launched in 2016 by the National Payments Corporation of India (NPCI). UPI has redefined financial transactions by enabling instant, real-time inter-bank transfers (Mistry, 2023). As of August 2023, UPI boasts the participation of 484 banks and has facilitated a staggering 10 billion transactions, totaling Rs 18,22,949.42 Crore(189 billion USD) (npc, 2023a). In India, 84 % of all digital transactions in 2022, amounting to 1.79 trillion dollars, were channeled through UPI (npc, 2023b). UPI’s rapid proliferation and high penetration in the Indian payment ecosystem underscore its transformative impact (Mistry, 2023).
@@
Despite UPI’s monumental success, a critical research gap exists regarding its influence on users’ spending habits. Our study addresses this crucial gap in existing research on UPI, specifically focusing on the impact of UPI on users’ spending habits. While previous studies and works (Sreenu, 2020) acknowledge UPI’s success in reshaping digital payments, they fall short of providing detailed insights into how this technology influences ”individual” spending behaviors. The existing literature often emphasizes macro-level implications and adoption rates. Multiple questions about changes in expenditure preferences, budgeting strategies, investment choices, long-term financial goals and overall financial behaviors remain unexplored, hindering a holistic understanding of UPI adoption’s intended and unintended consequences on users’ daily financial practices. This study seeks to fill these critical gaps by providing in-depth insights into the micro-level effects of UPI on spending behaviors. It adds valuable knowledge to the broader discourse on digital payment systems, stressing the importance of considering individual-level implications for a more holistic comprehension of financial technology’s transformative effects. Moreover, there’s a notable absence of research delving into potential variations in UPI’s impact across different demographic groups. Age, profession, and socio-economic status are overlooked in understanding how individuals interact with UPI and manage their finances.
We also examine users’ perceptions of security, investigating how trust in UPI’s transactional security affects their expenditure patterns. Additionally, the role of user trust in technology, particularly in UPI adoption, remains underexplored, prompting our study to unravel the intricate dynamics between trust, technology adoption, and financial decision-making. Recognizing the transformative nature of UPI, we also investigate the relationship between users’ financial literacy levels and their utilization of UPI, shedding light on how understanding financial principles shapes spending choices. Furthermore, our research extends its focus to entrepreneurs and businesses, probing into the effects of UPI adoption on their financial practices and business models. Another unique contribution of our research is analyzing UPI transactions’ integration with budgeting applications, employing a competitive lens to understand its impact on users’ awareness of spending patterns and financial discipline. Lastly, our study explores the dynamic nature of user satisfaction with UPI, considering system updates, security enhancements, and evolving user experiences. Collectively, our research endeavors aim to understand UPI’s influence on diverse aspects of users’ financial behaviors, contributing valuable insights to the evolving landscape of digital payments in India.
@@
Among the participants we surveyed, 74.2% of the participants confirmed increased spending due to UPI, with only 7% confirming reduced spending. Significantly, 91.5% of respondents reported satisfaction with their UPI usage. Also 95.2% of the users found making payments via UPI convenient. Our research also provides suggestions for UPI applications and various stakeholders to enhance digital payment systems, enabling users to make informed decisions and fostering responsible financial management. To the best of our knowledge, this is the first user study of UPI’s impact on individual spending behaviour in India.
@@ 2. Related Work
In recent times, there has been a growing interest in previously overlooked countries such as the BRIC nations, including Brazil, Russia, India, and China, where rapid economic and technological development is occurring (Kumar et al., 2011). Mobile payment systems are progressively influencing people’s socio-economic lives in South Asia (Asongu and Boateng, 2018; Asongu and Odhiambo, 2019; Ferreira and Perry, 2019; He et al., 2023; Lewis and Perry, 2019; Vashistha et al., 2019). India topped the list of the five countries, namely Brazil, China, Thailand, and South Korea, for digital payments by the end of 2022 with 89.5 million transactions (89., 2023). The sheer nature of mobile payments and their ubiquity, along with convenience, helps include individuals, eventually benefiting the country’s economy. Branchless banking (mobile money) apps are responsible for almost 30% of the GDP in various countries (ken, 2013; Ferreira and Perry, 2019).
@@
Historically, India has been characterized by high cash usage, with a substantial portion of the workforce engaged in the informal sector relying on cash transactions (Ghatak, 2017; Pal et al., 2018a). However, the government’s ”Digital India” vision, coupled with the rapid expansion of internet access, has propelled the adoption of digital payments. This transformation, from a cash-driven economy to a significant shift towards digital payments is also driven by various significant events like Demonetization and the COVID-19 pandemic. The rise of digital wallets and UPI platforms, including BHIM UPI (bhi, 2016), Paytm (pay, 2010), PhonePe (pho, 2015), and Google Pay (goo, 2012), etc. have played a pivotal role in reshaping the country’s financial landscape. Internet penetration in India increased from 14% in 2014 to over 46% in 2021, primarily attributed to affordable data tariffs introduced by regional mobile operators in 2016, with prices as low as Rs 15 (0.21 USD) per gigabyte (GB) of mobile data. As a result, India ascended to the second-largest global user base of the internet by 2018 (2ND, 2019; Jio, 2019).
Key events, such as Demonetization in 2016, in which, according to the government’s decision, most of the nation’s banknotes were declared out of legal tender aimed to combat corruption and terrorism (ter, 2017; Pal et al., 2018b; Kameswaran and Hulikal Muralidhar, 2019), further accelerated the shift towards digital payments. The COVID-19 pandemic acted as another catalyst, prompting the Reserve Bank of India to advocate for reduced cash usage to minimize physical contact, significantly increasing digital transactions. According to the World Bank, over 80 million Indian adults made their first digital merchant payment after the pandemic’s onset, further solidifying India’s position as a global digital payment leader (wor, 2022). Studies have indicated the positive impact of electronic transactions on trade and economic development in India (Abr, 2004; muk, 2011). The benefits include transparency, accountability, and reduced cash-related fraud, contributing to overall economic growth (Shiva, 2017).
@@
Despite progress, challenges remain, as only 35% of Indian adults use bank accounts for digital payments (glo, 2021). The low percentage (approx 5% of the Indian population) of income tax contributors emphasizes the potential role of digital payments in increasing tax collection, as highlighted by previous research (bax, 2005). The extensive penetration of smartphones in India, with 1.2 billion active mobile phone users, underscores the technological underpinnings of this digital transition (MIN, 2022). In 2016, a study by Goel (goe, 2016) found that technological advancements have played a significant role in empowering and introducing innovations in E-payment systems. Promoting digital payments is also associated with reducing cash-related crimes, as individuals tend to hold less physical cash, thus discouraging robbers from pursuing such crimes.
Research in the field has extensively explored the impact of mobile payment systems on economies and societies, focusing on their transformative effects. Notably, people from different countries and different strata of society manage finances differently (Fran Bennett, 2009; Kaye et al., 2014; Vines et al., 2011). This motivated us to understand the possible changes in their spending due to UPI and how Indian users keep track of their finances. Over the past decade, these systems have helped raise the standard of living and have revolutionized how money is used in developing economies.
@@
While previous research has broadly addressed the impact of mobile payment systems, our study focuses specifically on the financial behaviors of Indian users, particularly those using UPI. While the existing literature has focused on macro-level implications, multiple questions about changes in expenditure preferences, budgeting strategies, long-term financial goals and overall financial behaviors remain unexplored. This study seeks to fill these critical gaps by providing in-depth insights into the micro-level effects of UPI on spending behaviors, thus, giving a holistic understanding of UPI adoption’s intended and unintended consequences on users’ daily financial practices. Additionally, our research pays particular attention to the concerns and preferences of users, shedding light on factors that may influence the widespread acceptance of mobile payment systems in the region. Moreover, our work also aims to understand user trust in UPI technology. The study aims to uncover the factors that encourage users to trust the UPI technology, considering the backdrop of rising scams (sca, 2021, 2023). By delving into this aspect, we aim to provide insights beyond transaction numbers, contributing to a more comprehensive understanding of India’s evolving digital payments landscape.
@@
Evidence suggests that individuals often under-save, fail to invest wisely, and are frequently indebted (stu, 2012). Increasing financial literacy improves financial decision-making, leading governments worldwide to recognize it as a critical life skill, launching financial education initiatives to help young people acquire this skill (asi, 2014; rbi, 2020). Multiple studies have highlighted the importance of financial inclusion and literacy, with one study pointing out how roadblocks to digital or financial literacy impede digital financial inclusion (Srivastava, 2022). A study by Agarwalla et al. (iim, 2014) focused on Indian working youth, stating that focused intervention is needed to improve the ”financial attitude” of men and the ”financial behavior” of women. French et al. (French et al., 2021) examined the effectiveness of personal finance smartphone apps in enhancing financial capability for low-income individuals by developing four specific apps. While cashless transactions promote technology literacy (tan, 2019), there is a visible gap in understanding how or if it promotes financial literacy. A financial activity study was carried out in 2014, tracking people’s transactional activity over time and interviewing them about their practices, understandings, needs, concerns, and expectations of present and future financial technologies to explore how people might interact with such digital payment systems (Lewis and Perry, 2019). As demonstrated by Lusardi et al. (llm, 2012), financial literacy training at an early stage in the working career significantly impacts financial outcomes due to compounded returns over a more extended period. Several studies worldwide indicate that a significant fraction of individuals have a low level of financial knowledge, associated with sub-optimal financial outcomes in various areas such as retirement planning (Lusardi and Mitchell, 2014; Lus, 2009; roo, 2011). In contrast to the existing studies, our study aims to explore how financial literacy impacts spending habits, especially with the onset of cashless transactions (UPI) in the context of India.
Finally, in this study, we conduct a comparative analysis of budgeting apps in India, an in-depth analysis of customer expectations from financial management or budgeting apps, and an understanding of specific user preferences that can act as a baseline for concerned developers, authorities, and stakeholders.
@@ 3. Methodology
3.1. RESEARCH DESIGN - MIXED DESIGN
In our research design, we employed a mixed methods (mul, [n. d.]) approach to comprehensively investigate the impact of India’s Unified Payments Interface (UPI) on individual spending behaviors and financial practices. This approach allowed us to gain a holistic understanding of how UPI has influenced users’ spending habits. We utilized two primary data collection techniques: surveys and interviews.
We employed a random sampling method for the survey component, targeting individuals aged 18 and above who may/may not have used UPI for payments. Our survey was conducted through Google Forms to ensure easy accessibility and respondent anonymity (by not asking for their personal information like name, phone number, etc as seen in Section 7). It garnered 276 responses from a diverse participant pool spanning various age groups and occupations, including working professionals, students, and business owners. It was floated across various social media platforms and groups and no questions were marked as mandatory, however, respondents were encouraged to provide honest and accurate responses as much as possible.
To ensure the quality and relevance of the responses, we implemented stringent criteria for validity. Frivolous responses containing irrelevant information were excluded, as were responses that deviated from the expected format. For instance, in the question regarding the initiation of UPI usage (”When did you start using UPI?”), any response indicating usage before 2016 was deemed invalid, considering that UPI was publicly launched that year. This validation process resulted in 235 valid responses for inclusion in our research work. Furthermore, to maintain consistency, we addressed ambiguous responses to the question about the commencement of UPI usage. For instance, answers like ”2 years back” were standardized to 2021, aligning with the survey release date of 2023. By adopting these rigorous validation measures, we ensured a reliable and high-quality dataset for our analysis.
In addition to surveys, we conducted in-depth interviews to dive deeper into the experiences and perspectives of UPI users. We randomly selected interviewees from our survey respondents who expressed an interest in participating in discussions. We conducted 20 (13 Male/ 7 Female) interviews over Google Meet and the transcripts of these interviews were further analysed. The interviewees provided verbal consent for the interview.
3.2. ETHICAL CONSIDERATION
@@
3.3. LIMITATION
One notable constraint of this study is the geographic scope of our data collection efforts. The research exclusively focused on respondents from urban areas in India, where technology adoption and usage, including electronic payment methods like UPI, tend to be higher and well integrated into daily life. Urban living typically involves residing in densely populated areas with access to advanced infrastructure and services, while rural living often means residing in less densely populated areas with a focus on agriculture and simpler infrastructure. As a result, the findings may not encompass the diverse spending habits of rural areas. They may not accurately represent the impact of UPI on the spending habits of living in rural areas in India.
@@ 4. Evaluations
4.1. QUANTITATIVE ANALYSIS ( ANALYSIS OF SURVEY RESULTS)
@@
Our survey received 235 valid responses from a diverse range of respondents from various demographics and backgrounds. Among the respondents, 51.6% identified as students, 42% as working professionals, and 4.7% as business owners. Regarding gender distribution, the data shows that 77% of respondents were male, 19.7% were female, and 3.3% preferred not to disclose their gender. Regarding the adoption timeline of UPI as shown in figure 1(a), a notable 12% of respondents started using it in 2016 after the ’Demonetization’ of Rs 500 (6.03 USD) and Rs 1000 (12.07 USD) notes to combat black money and counterfeiting. Followed by 7.60% in 2017, 13.60% in 2018, 15.60% in 2019, 16% in 2020, 17.6% in 2021, 11.20% in 2022, and 2.40% in 2023. This diverse range of adoption dates reflects users’ varying entry points into the UPI ecosystem, often aligning with significant events like demonetization and pandemic-related lockdowns.
@@
To understand the impact of UPI adoption, 74.2% of the survey respondents admitted that UPI had led to increased spending, while only 7% said that it had decreased spending. Additionally, 18.8% of the survey respondents felt no changes in their spending habits after adopting UPI. Regarding satisfaction levels, as shown in figure 2(a), 55% of the survey respondents indicated they were highly satisfied with their overall UPI experience. Additionally, 36.5% reported satisfaction, and 7.7% remained neutral in their assessment. Therefore we can conclude that 91.5% (55% + 36.5%) of the respondents are satisfied with their overall experience using UPI for payments. When asked to rate their convenience with UPI for making payments on a scale from 1 (lowest) to 5 (highest), as shown in figure 4, the survey results indicate that 69.50% of respondents rated ”5” (Highly convenient), 25.7% rated ”4” (convenient), and 3.7% were ”3” (Neutral). Therefore we can conclude that 95.2% (69.5% + 25.7%) of the respondents find using UPI for payments convenient. Notably, only 2-3 respondents reported dissatisfaction with UPI as a payment method, indicating that most were satisfied with their experience with UPI usage.
@@
Regarding daily spending using UPI, respondents reported a wide range, with amounts varying from Rs 50 (0.60 USD) to 1 lakh (1,202.66 USD). On average, 15.7% of respondents mentioned their daily average spending is around Rs 200 (2.41 USD). When respondents were asked about their frequency of using UPI for payments, as shown in figure 1(b), an overwhelming 81.1% indicated that they use it daily. Additionally, 15.6% reported using UPI weekly, while 1.1% said they use it monthly, 1.90% mentioned using it occasionally, and 0.30% mentioned rarely. This data underscores the widespread and frequent usage of UPI among the surveyed individuals.
@@
When asked about exceeding their budget or overspending due to UPI usage, as shown in the figure 3(a), 59.8% of respondents acknowledged experiencing such situations, while 39.8% believed they had not. Additionally, respondents had diverse opinions when asked about the impact of using UPI on their budget adherence, as shown in figure 2(b); approximately 23% strongly agreed (rating it as 5), 23.3% agreed (rating it as 4), 26.7% remained neutral (rating it as 3), 15.6% disagreed (rating it as 2), and 11.5% reported strongly disagree (rating it as 1), indicating that UPI has had an impact on their ability to stick to their budget and control overspending.
Figure 1. (a) When did users start using UPI? (b) How often do users use UPI?
Figure 2. (a) Satisfaction Level with UPI (b) UPI Impact on Budget
Figure 3. (a) Budget exceedance due to UPI (b) How often do people keep track of transaction history?
@@
Respondents employed various methods to manage their expenses, including popular platforms like Splitwise (spl, 2011) and Paytm (pay, 2010), bank statements, and applications such as SBI YONO (yon, 2017) and Samsung Notes (sno, 2016). Other tools included Excel sheets and dedicated finance apps like MyMoneyy (mym, 2022), Honeybook (hon, 2013), Monefy (my, 2016), Payhawk (pay, 2018) and Money Manager (mm, 2020). Notably, many respondents mentioned that they need to check their messages or transaction history regularly. This observation sheds light on users’ varied financial management, literacy, and discipline approaches. Furthermore, participants provided additional comments, and some expressed their willingness to participate in interviews. These findings illustrate a wide range of responses and opinions among UPI users, with a significant proportion acknowledging increased spending and high overall satisfaction with the platform and its convenience.
4.2. QUALITATIVE ANALYSIS (ANALYSIS OF INTERVIEW RESULTS)

### v2

@@ Abstract.
The emergence of digital payment systems has transformed how individuals conduct financial transactions, offering convenience, security, and efficiency. One groundbreaking innovation making waves in the Indian financial landscape is the Unified Payments Interface (UPI). Existing work has explored how digital payments benefit a country’s economy and GDP. However, our study explores how the introduction of UPI has influenced spending behavior among Indian users on an ”individual” level. We gathered 235 valid survey responses encompassing diverse demographics and interviewed 20 survey respondents. Approximately 75% of the survey respondents reported increased spending due to UPI, with only 7% indicating reduced spending. Significantly, 91.5% of the respondents reported satisfaction with their UPI usage. Also, 95.2% of the survey respondents found making payments via UPI convenient. Our research also provides suggestions for UPI applications and various stakeholders to enhance digital payment systems, enabling users to make informed decisions and fostering responsible financial management. *Equal Contribution
@@
journalyear: 2024copyright: rightsretainedconference: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (CHI EA ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613905.3651050isbn: 979-8-4007-0331-7/24/05ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
An E-payment is generally understood as an economic situation where the buying and selling of goods are handled without physical cash, often facilitated through E-payment or checks (Pau, 2012). Digital payments have become integral to modern financial landscapes, offering swift, secure, and convenient electronic transactions that transcend traditional methods. This global shift towards digital transactions has been accelerated by the widespread adoption of internet and mobile technologies, reshaping financial practices across the world (Alkhowaiter, 2020). In 2022 alone, the volume of real-time digital transactions reached an unprecedented 195 million (89., 2023).
@@
India, in particular, has emerged as a global leader in digital transactions, accounting for 46 % of the total global digital transactions in 2022 (89., 2023). Unified Payments Interface (UPI), launched in 2016 by the National Payments Corporation of India (NPCI), has redefined financial transactions by enabling instant, real-time inter-bank transfers (Mistry, 2023). As of August 2023, UPI has facilitated a staggering 10 billion transactions, totaling Rs 18,22,949.42 Crore(189 billion USD) (npc, 2023a). In India, 84 % of all digital transactions in 2022, amounting to 1.79 trillion dollars, were channeled through UPI (npc, 2023b). UPI’s rapid proliferation and high penetration in the Indian payment ecosystem underscore its transformative impact (Mistry, 2023).
@@
Despite UPI’s monumental success, a critical research gap exists regarding its influence on users’ spending habits. Our study addresses this crucial gap in existing research on UPI, specifically focusing on the impact of UPI on users’ spending habits. While previous studies and works (Sreenu, 2020) acknowledge UPI’s success in reshaping digital payments, they fall short of providing detailed insights into how this technology influences ”individual” spending behaviors. The existing literature often emphasizes macro-level implications and adoption rates. Multiple questions about changes in expenditure preferences, budgeting strategies, investment choices, long-term financial goals and overall financial behaviors remain unexplored, hindering a holistic understanding of UPI adoption’s intended and unintended consequences on users’ daily financial practices. This study seeks to fill these critical gaps by providing in-depth insights into the micro-level effects of UPI on spending behaviors. Moreover, there’s a notable absence of research delving into potential variations in UPI’s impact across different demographic groups like age, profession, and socio-economic status, etc.
We explore users’ perceptions of security and how trust in UPI influences expenditure patterns, unraveling the intricate dynamics between trust, technology adoption, and financial decision-making. Additionally, we investigate the relationship between users’ financial literacy levels and UPI utilization, shedding light on how understanding financial principles shapes spending choices. Our research also analyzes the existing budgeting applications, employing a competitive lens to understand its impact on users’ awareness of spending patterns and financial discipline.
@@
Among the participants we surveyed, 74.2% of the survey participants confirmed increased spending due to UPI, with only 7% confirming reduced spending. Significantly, 91.5% of the survey respondents reported satisfaction with their UPI usage and 95.2% of the survey respondents found making payments via UPI convenient. Our research also provides suggestions for UPI applications and various stakeholders to enhance digital payment systems, enabling users to make informed decisions and fostering responsible financial management. To the best of our knowledge, this is the first user study of UPI’s impact on individual spending behaviour in India.
@@ 2. Related Work
Mobile payment systems are progressively influencing people’s socio-economic lives in South Asia (Asongu and Boateng, 2018; Asongu and Odhiambo, 2019; Ferreira and Perry, 2019; He et al., 2023; Lewis and Perry, 2019; Vashistha et al., 2019). India topped the list of the five countries, namely Brazil, China, Thailand, and South Korea, for digital payments by the end of 2022 with 89.5 million transactions (89., 2023). The sheer nature of mobile payments and their ubiquity, along with convenience, helps include individuals, eventually benefiting the country’s economy (ken, 2013; Ferreira and Perry, 2019).
@@
Historically, India has been characterized by high cash usage, with a substantial portion of the workforce engaged in the informal sector relying on cash transactions (Ghatak, 2017; Pal et al., 2018a). The transformation from a cash-driven economy to a significant shift towards digital payments is also driven by various significant events like government’s ”Digital India” vision, Demonetization (ter, 2017; Pal et al., 2018b; Kameswaran and Hulikal Muralidhar, 2019) and COVID-19 pandemic. According to the World Bank, over 80 million Indian adults made their first digital merchant payment after the pandemic’s onset(wor, 2022). Studies have indicated the positive impact of electronic transactions on trade and economic development in India (Abr, 2004; muk, 2011). The benefits include transparency, accountability, and reduced cash-related fraud, contributing to overall economic growth (Shiva, 2017). The surge of digital wallets and UPI platforms, including BHIM UPI (bhi, 2016), Paytm (pay, 2010), PhonePe (pho, 2015), and Google Pay (goo, 2012), has been instrumental in reshaping India’s financial landscape. Internet penetration in the country rose from 14% in 2014 to over 46% in 2021, driven by affordable data tariffs introduced by regional mobile operators in 2016. This contributed to India becoming the second-largest global user base of the internet by 2018 (2ND, 2019; Jio, 2019).
@@
Despite progress, challenges remain, as only 35% of Indian adults use bank accounts for digital payments (glo, 2021). The low percentage (approx 5% of the Indian population) of income tax contributors emphasizes the potential role of digital payments in increasing tax collection, as highlighted by previous research (bax, 2005). The extensive penetration of smartphones in India, with 1.2 billion active mobile phone users, underscores the technological underpinnings of this digital transition (MIN, 2022). In 2016, a study by Goel (goe, 2016) found that technological advancements have played a significant role in empowering and introducing innovations in E-payment systems.
@@
Research in the field has extensively explored the impact of mobile payment systems on economies and societies, focusing on their transformative effects. Notably, people from different countries and different strata of society manage finances differently (Fran Bennett, 2009; Kaye et al., 2014; Vines et al., 2011). This motivated us to understand the possible changes in their spending due to UPI and how Indian users keep track of their finances. This study addresses the micro-level effects of UPI on spending behaviors, filling gaps in existing literature focused on macro-level implications. It explores changes in expenditure preferences, budgeting strategies, and long-term financial goals. The research aims to comprehensively understand the consequences of UPI adoption on users’ spending behavior while highlighting user concerns and preferences that influence the widespread acceptance of mobile payment systems in the region. The study also aims to provide insights beyond transaction numbers and uncover the factors that encourage users to trust the UPI technology, considering the backdrop of rising scams (sca, 2021, 2023).
@@
Evidence suggests that individuals often under-save, fail to invest wisely, and are frequently indebted (stu, 2012). Several studies indicate that a significant fraction of individuals has a low level of financial knowledge, associated with sub-optimal financial outcomes in various areas such as retirement planning (Lusardi and Mitchell, 2014; Lus, 2009; roo, 2011; llm, 2012). Increasing financial literacy improves financial decision-making, leading governments worldwide to recognize it as a critical life skill, launching financial education initiatives to help young people acquire this skill (asi, 2014; rbi, 2020). Multiple studies have highlighted the importance of financial inclusion and literacy, with one study pointing out how roadblocks to digital or financial literacy impede digital financial inclusion (Srivastava, 2022). While cashless transactions promote technology literacy (tan, 2019), there is a visible gap in understanding how or if it promotes financial literacy. In contrast to the existing studies, our study aims to explore how financial literacy impacts spending habits, especially with the onset of cashless transactions (UPI) in the context of India.
@@ 3. Methodology
3.1. RESEARCH DESIGN
We employed a mixed methods (mul, [n. d.]) approach to investigate the impact of UPI on individual spending behaviors and financial practices. We utilized two primary data collection techniques: surveys and interviews. We used a random sampling method for the survey component, targeting individuals aged 18 and above. Our survey was conducted through Google Forms to ensure easy accessibility and respondent anonymity (by not asking for their personal information like name, phone number, etc). It garnered 276 responses from a diverse participant pool spanning various age groups and occupations. To ensure the quality, consistency and relevance of the responses, we implemented stringent criteria for validity. Frivolous responses containing irrelevant information were excluded, for instance, in the question regarding the initiation of UPI usage, any response indicating usage before 2016 was deemed invalid, considering that UPI was publicly launched that year. This validation process resulted in 235 valid responses. We conducted in-depth semi-structured interviews by randomly selecting interviewees from our survey respondents who expressed an interest in participating. We conducted 20 (13 male & 7 female) interviews and they provided verbal consent for the same. Survey and interview questions are presented in Appendix A and Appendix B respectively.
3.2. ETHICAL CONSIDERATIONS
@@
3.3. LIMITATIONS
One notable constraint of this study is the geographic scope of our data collection efforts. The research exclusively focused on respondents from urban areas in India, where technology adoption and usage, including electronic payment methods like UPI, tend to be higher and well integrated into daily life. As a result, the findings may not encompass the diverse spending habits. They may not accurately represent the impact of UPI on the spending habits of those living in rural areas in India.
@@ 4. Evaluations
4.1. QUANTITATIVE ANALYSIS OF SURVEY RESPONSES
@@
Our survey received 235 valid responses from a diverse range of respondents from various demographics and backgrounds. Among the respondents, 51.6% identified as students, 42% as working professionals, and 4.7% as business owners. Regarding gender distribution, the data shows that 77% of respondents were male, 19.7% were female, and 3.3% preferred not to disclose their gender. Regarding the adoption timeline of UPI (figure 1), we can conclude that the diverse range of adoption dates reflects users’ varying entry points into the UPI ecosystem, often aligning with significant events like demonetization and pandemic-related lockdowns.
@@
To understand the impact of UPI adoption, 74.2% of the survey respondents admitted that UPI had led to increased spending, while only 7% said that it had decreased spending. Additionally, 18.8% of the survey respondents felt no changes in their spending habits after adopting UPI. Regarding satisfaction levels (figure 3), 55% of the survey respondents indicated they were highly satisfied with their overall UPI experience. Additionally, 36.5% reported satisfaction, and 7.7% remained neutral in their assessment. Therefore we can conclude that 91.5% (55% + 36.5%) of the respondents are satisfied with their overall experience using UPI for payments. When asked to rate their convenience with UPI for making payments (figure 7), the survey results indicate that 69.50% of respondents rated ”Highly convenient”, 25.7% rated ”Convenient, and 3.7% were ”Neutral” on the same. Therefore we can conclude that 95.2% (69.5% + 25.7%) of the respondents find using UPI for payments convenient. Notably, only 2-3 respondents reported dissatisfaction with UPI as a payment method, indicating that most were satisfied with their experience.
@@
Respondents reported a varied daily spending range using UPI, from Rs 50 (0.60 USD) to 1 lakh (1,202.66 USD). On average, 15.7% mentioned a daily spending of around Rs 200 (2.41 USD). Regarding the frequency of UPI usage for payments (figure 2), a significant 81.1% use it daily, 15.6% use it weekly, 1.1% use it monthly, 1.90% use it occasionally, and 0.30% use it rarely. This data highlights the widespread and frequent adoption of UPI among surveyed individuals.
@@
When asked about overspending due to UPI usage (figure 5), 59.8% of respondents acknowledged experiencing such situations, while 39.8% believed they had not. Additionally, respondents had diverse opinions when asked about the impact of using UPI on their budget adherence (figure 4); approximately 23% strongly agreed, 23.3% agreed, 26.7% remained neutral, 15.6% disagreed, and 11.5% reported strongly disagree, indicating that UPI has had an impact on their ability to stick to their budget and control overspending.
Respondents employed various methods to manage their expenses, like checking their transaction history, Splitwise (spl, 2011) SBI YONO (yon, 2017), Samsung Notes (sno, 2016), using excel sheets and dedicated finance apps like MyMoneyy (mym, 2022), Honeybook (hon, 2013), Monefy (my, 2016), Payhawk (pay, 2018) and Money Manager (mm, 2020). This observation highlights diverse financial management approaches among UPI users, showcasing varied responses and opinions. A significant proportion acknowledges increased spending along with high overall satisfaction with UPI and its convenience. Our comprehensive analysis of various financial tracking and budget management apps in the Indian market is summarized in Table 1 in Section 5.
Figure 1. When did users start using UPI?
Figure 2. How often do users use UPI?
Figure 3. Satisfaction Level with UPI
Figure 4. UPI Impact on Budget
Figure 5. Budget exceedance due to UPI
Figure 6. How often do people keep track of transaction history?
@@
4.2. QUALITATIVE ANALYSIS OF INTERVIEWS